# Federated Learning for Traffic Flow Prediction with Synthetic Data Augmentation

### v1

@@ Abstract.
Deep-learning based traffic prediction models require vast amounts of data to learn embedded spatial and temporal dependencies. The inherent privacy and commercial sensitivity of such data has encouraged a shift towards decentralised data-driven methods, such as Federated Learning (FL). Under a traditional Machine Learning paradigm, traffic flow prediction models can capture spatial and temporal relationships within centralised data. In reality, traffic data is likely distributed across separate data silos owned by multiple stakeholders. In this work, a cross-silo FL setting is motivated to facilitate stakeholder collaboration for optimal traffic flow prediction applications. This work introduces an FL framework, referred to as FedTPS, to generate synthetic data to augment each client’s local dataset by training a diffusion-based trajectory generation model through FL. The proposed framework is evaluated on a large-scale real world ride-sharing dataset using various FL methods and Traffic Flow Prediction models, including a novel prediction model we introduce, which leverages Temporal and Graph Attention mechanisms to learn the Spatio-Temporal dependencies embedded within regional traffic flow data. Experimental results show that FedTPS outperforms multiple other FL baselines with respect to global model performance.
@@
Federated Learning, Traffic Flow Prediction, Synthetic Data Generation
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; December, 2024; ccs: Information systems Data streamingccs: Information systems Spatial-temporal systems
1. Introduction
@@
Traffic prediction is an integral aspect of modern Intelligent Transport Systems (ITSs) with innate implications on the safety and productivity of inhabitants; it facilitates the optimisation of traffic management systems, and the dissemination of pertinent information for commuters. This in turn has profound implications on social, environmental, and economic factors. Deep-learning based models have demonstrated exceptional accuracy in their ability to learn spatio-temporal relationships from traffic data and predict future traffic states (Yu et al., 2018; Li et al., 2018; Wu et al., 2019; Tan et al., 2023; Cho et al., 2014).
@@
The majority of previous works towards spatio-temporal prediction aim to train a model from a centralised perspective where the collective data is available one a single device. Centralising city-wide ITS data in real-world settings is a difficult task; ITSs represent a partnership between transport stakeholders such as local government authorities, ride-sharing companies, and micro-mobility organisations. Sharing journey data has implications on user privacy (GDPR), as well as commercial implications. Therefore, developing frameworks for applicable decentralised data-driven methods is of paramount importance within the ITS sector to enable stakeholder collaboration for the development of traffic prediction models. Federated Learning (FL) (McMahan et al., 2017) provides a solution for training optimal traffic prediction models collaboratively by decentralising the training process of a traffic prediction model. Although recent works in FL demonstrate its capacity to train a global model which learns spatio-temporal dependencies across distributed data, these works do not directly address the issue of data heterogeneity (Hu et al., 2023), or instead focus on the client selection strategy or parameter aggregation function (Meng et al., 2021; Yuan et al., 2023; Liu et al., 2020; Zhang et al., 2021b, a, 2022; Qi et al., 2023).
@@
To tackle the issue of data heterogeneity and imbalanced data partitions across clients, data sharing mechanisms for FL applications in highly heterogeneous settings were explored in (Zhao et al., 2018) wherein a small partition of real data is distributed to clients to augment their local datasets in a cross-device setting. Sharing real data with clients introduces a compromise between accuracy and privacy of data. Recent works have looked at training generative models within an FL framework for image classification tasks, to augment local client datasets without sharing data that would compromise data privacy (Li et al., 2023; Triastcyn and Faltings, 2020; Behera et al., 2022). For applications beyond image classification where data contains spatial and temporal dependencies and heterogeneity is principally due to quantity and feature-based skews, limited work exists on the benefits of augmenting FL client datasets using synthetic data.
@@
Emerging research in the field of generative models for trajectory generation have demonstrated the exceptional capacity for diffusion-based models to generate synthetic data that captures spatio-temporal cross-dependencies of traffic data while maintaining the privacy of real data (Zhu et al., 2023). Leveraging recent works in FL as well as in the field of generative models for spatio-temporal data generation, we present a framework for Federated Traffic Prediction with Synthetic data augmentation (FedTPS). This framework first trains a diffusion-based model for traffic data generation through FL. This model is used to generate a synthetic dataset developed from the learned global data distribution. A traffic flow prediction model is then trained through FL, wherein each client’s local dataset has been augmented with the generated synthetic dataset.
@@
In real-world implementations of traffic flow prediction models with data silos, the data owned by separate clients may contain inherent skews, contributing to data heterogeneity across client datasets. The proposed FedTPS framework directly inhibits the negative effects of data heterogeneity when training a global through FL, while also increasing the amount of training data in each client’s local dataset. To the best of our knowledge, this work is the first to explore the benefits of exploiting synthetic data for FL in traffic prediction applications.
@@
  • We assess the performance of several traffic flow prediction models trained under an FL setting. We also develop a novel traffic prediction model, the Graph Attention and Temporal Attention Unit (GATAU), which builds upon the existing spatio-temporal model, the Temporal Attention Unit (TAU) (Tan et al., 2023), to also leverage Graph Attention mechanisms.
@@
  • We propose a novel FL framework, FedTPS, which trains a federated generative model to produce a synthetic dataset. This synthetic data is used to augment client datasets for traffic flow prediction applications, and improve the predictive performance of the FL model.
@@
  • Our experiments are validated on a real-world dataset to analyse the impact of data augmentation on traffic flow prediction model performance, and demonstrate improved results relative to multiple other FL frameworks designed to address data heterogeneity across data-silos.
2. Related Works
2.1. Federated Learning
@@
The distributed nature of the data in a FL setting incurs a detrimental effect on model accuracy relative to centralised training paradigms due to the fact that the data distribution among clients is non-Independent and Identically Distributed (non-IID). In cross-silo FL settings, each client may represent a data center which stores large amounts of data aggregated from many devices, and owned by a particular organisation. FL has been extended to optimise the training of the global model (Reddi et al., 2021), as well as to improve its applicability in non-IID settings (Li et al., 2020). Yet such methods focus on optimising the FL framework rather than address the data heterogeneity issues directly.
@@
Existing data augmentation approaches for FL focus on image classification tasks, and typically rely on Generative Adversarial Networks (GANs) (Triastcyn and Faltings, 2020; Li et al., 2023; Behera et al., 2022), or on sharing mean latent representations of client data (Yoon et al., 2021; Chen and Vikalo, 2023). Unlike image generation applications, traffic data generation requires modelling the dynamics of data over both spatial and temporal dimensions. Diferentially-private representations of spatio-temporal data can be produced by altering real data (Armstrong et al., 1999; Zandbergen, 2014), or mixing partial aspects of the data (Nergiz et al., 2008). These processes affect the spatio-temporal characteristics of the data itself, however. Recent work into diffusion-based models for spatio-temporal data generation demonstrate their capacity to generate high-fidelity synthetic trajectories which capture the spatio-temporal characteristics of sample traffic data, while ensuring the privacy of the real data.
DiffTraj (Zhu et al., 2023) is a conditional diffusion-based trajectory generation model which leverages Traj-UNet, a network based on the UNet architecture (Ronneberger et al., 2015) as well as an intermediary attention-based transition module, to model the noise transitions at each time step in the diffusion model processes. The model embeds additional context, such as the starting region and departure time of a trip, as conditional information which is implanted by a Wide and Deep network (Cheng et al., 2016) to enhance the generative capacity of the model. To avoid memorisation of trajectories based on conditional information, which would implicate the privacy of real data, DiffTraj uses the classifier-free diffusion guidance method (Ho and Salimans, 2021) which jointly trains an unconditional model to increase the diversity of generated trajectories and avoid deterministic data generation.
2.2. Traffic Flow Prediction
To extract temporal dependencies, deep-learning based models typically employ either Recurrent Neural Networks (RNNs) based architectures (Li et al., 2018; Liu et al., 2020), temporal Convolutional Neural Networks (CNNs) which convolve around the time dimension of traffic flow data (Yu et al., 2018; Wu et al., 2019), or temporal Attention mechanisms to capture the evolution of the data around the time dimension (Tan et al., 2023).
Spatial relationships embedded within data can be captured through CNNs which convolve around the spatial dimension of data (Tan et al., 2023). Traffic data can also be represented as graph networks, enabling Graph Convolutional Networks (GCNs) to capture spatial relationships within the data. GCN operations are dependent on the graph structure, which demands expert contextual knowledge to adequately capture inter-region relationships. Graph Attention Networks (GATs) (Veličković et al., 2018) implement a self-attention mechanism to learn the relevance between a given node and its neighbouring nodes, and hence are independent of the graph structure. To capture spatio-temporal dependencies within data, models such as STGCN (Yu et al., 2018) implement GCNs with temporal CNNs to capture spatio-temporal dependencies within data, while DCRNN (Li et al., 2018) leverages GCNs with RNN-based models. GraphWaveNet (Wu et al., 2019) also utilises temporal CNNs and GCNs, however it learns an adjacency matrix for the graph structured data. The Temporal Attention Unit (TAU) (Tan et al., 2023) utilises CNNs and temporal Attention mechanisms to capture both spatial and temporal relationships within data.
@@ 2.2. Traffic Flow Prediction
Existing work on FL applied for traffic prediction implement novel client selection strategies to cluster clients with similar spatial characteristics (Zhang et al., 2022, 2021a; Meng et al., 2021). Other such works look to adapt the aggregation procedure of the FL setting (Zhang et al., 2021b; Yuan et al., 2023; Qi et al., 2023). FedTDP (Hu et al., 2023) explores the data heterogeneity issue which may be found in real-world traffic data silos, yet does not propose a method for dealing with the inherent data heterogeneity across client data silos.
3. Problem Formulation
Figure 1. Satellite image of Chengdu, showing a sample of trajectories shown in cyan representing $X_{c}$, and where traffic flow into the red regions represents $Q_{c}$.
@@ 3. Problem Formulation
Chengdu map with trajectories and regional inflow
Definition 1 (Data Generation): Trajectory data can be defined as a sequence $x$ of GPS latitude-longitude pairs. The objective of data generation is, given a set of GPS trajectories, $\mathcal{X}=\{\bm{x}^{1},\bm{x}^{2},\ldots,\bm{x}^{m}\}$, where $\bm{x}^{i}$ is a sequence of GPS latitude-longitude pairs, learn a model, G, that can generate a set of synthetic trajectories, $,X_{\textnormal{synth}}=\{\bm{x_{\textnormal{synth}}}^{1},\newline \bm{x_{\textnormal{synth}}}^{2},$$\ldots,\bm{x_{\textnormal{synth}}}^{m}\}$, where $\bm{x_{\textnormal{synth}}}^{i}$ is a sequence of synthetic latitude-longitude pairs, and where the synthetic trajectories retain the real data’s spatial and temporal characteristics without revealing information pertaining to the real trajectories. To emulate the setting of various organisations each with ownership over separate data silos, the global trajectory dataset, $X_{\textnormal{glob}}$, is partitioned into $C$ clients such that $X_{\textnormal{glob}}=(X_{1}\cup X_{2}\cup...\cup X_{C})$, where $X_{c}$ is the trajectory dataset owned by client $c$. The FL task is to train model G through FL using the distributed client datasets, and develop a synthetic dataset $X_{\textnormal{synth}}$.
Definition 2 (Traffic Flow Prediction): A city can be discretised into $N$ disjoint regions, and $q_{t}\in\mathbb{R}^{N}$ can be defined as the traffic inflow for all regions $N$ at time step $t$. Given historical traffic flow $\bm{Q}=(q_{1},q_{2},...,q_{T})\in\mathbb{R}^{NxT}$, train a model to predict the traffic inflow of all regions at future time step $T+\tau$, hence predict $q_{T+\tau}$. Based on the partitioned trajectory data $X_{\textnormal{glob}}$, client $c$ can derive its regional traffic inflow data $\bm{Q}_{c}$, and can also augment its own local dataset using the synthetic dataset $X_{\textnormal{synth}}$. The federated traffic flow prediction task is then to collaboratively train a global prediction model using the distributed client datasets with synthetic data augmentation.
4. Methodology
This work aims to address a real-world setting of multiple organisations, each with their independent dataset of vehicle trajectories, collaborating through FL to train a trajectory generation model and subsequent traffic flow prediction model. To emulate this scenario, the trajectory dataset is segmented into subsets to represent local client datasets in a cross-silo FL setting. Within this FL setting, a server acts as a secure, trusted central system which coordinates federated training between clients, and whose role is further detailed in Algorithms 1 and 2. The proposed FedTPS framework first aims to train a generative model through the partitioned vehicle trajectory datasets, and generate a synthetic dataset from the global distribution of data. The second step aims to train a traffic flow prediction model using each client’s traffic flow data, and augment each client’s data with the synthetic data generated from the first step. This will reduce the effects of data heterogeneity present across the client datasets, and increase the amount of training data available to clients.
4.1. Federated Generative Model
As a probabilistic generative model, a diffusion model aims to generate synthetic data by modeling two sequential processes: 1) A series of forward processes which gradually injects noise into the data, and 2) The reverse of these processes, which learn to capture the original data distribution from a noisier version (Sohl-Dickstein et al., 2015). Recent works have applied diffusion-based models for high-fidelity data generation in various domains (Gong et al., 2023; Ho et al., 2022; Xu et al., 2022). The forward process is defined as a Markov chain with $F$ Gaussian transitions that map the real data $x_{0}\sim g(x_{0})$ to $x_{F}$, a latent variable with the same dimensionality as $x_{0}$, according to
(1) $g(x_{1},...,x_{F}|x_{0})=\prod^{F}_{f=1}g(x_{f}|x_{f-1}),$
(2) $g(x_{f}|x_{f-1})=\mathcal{N}(x_{f};\sqrt{1-\beta_{f}}x_{f-1},\beta_{f}\mathbf{I}),$
(3) $p_{\theta}(x_{0},...,x_{f-1}|x_{F})=p(x_{F})\prod^{F}_{f=1}p_{\theta}(x_{f-1}|x_{f}),$
(4) $p_{\theta}(x_{f-1}|x_{f})=\mathcal{N}(x_{f-1};\mu_{\theta}(x_{f},f),\sigma_{\theta}(x_{f},f)^{2}\mathbf{I}),$
(5) $\mathcal{L}_{c}(\theta):=\mathbb{E}_{x_{0}\in X_{c},\epsilon,f}\left[||\epsilon-\epsilon_{\theta}(x_{f},f)||^{2}\right].$
Conditional diffusion models enable additional information to be embedded into the reverse denoising process as conditional variables into Eq. (4). The diffusion-based generative model DiffTraj (Zhu et al., 2023) is used as the conditional model for synthetic data generation in FedTPS, as illustrated in Algorithm 1.
4.1.1. Federated Framework
In a scenario where trajectory data is segmented across data silos from multiple organisations, a diffusion model cannot be trained under a traditional framework since the training data is not centralised. Given the vast amount of data required to train a conditional diffusion-based model such as DiffTraj, each client may not be able to individually train an accurate data generation model which captures the spatio-temporal characteristics across the city with high fidelity.
Figure 2. FedTPS Framework for training a federated generative model. $X^{\textnormal{CO}}_{c}$ represents client $c$’s conditional observations required for the DiffTraj model.
[…]
System level diagram of server and clients in FL setting for DiffTraj
Under a federated setting, the updated parameters from each client’s local training on the global model are communicated to a central server which then aggregates the results to update the global model. FedAvg aggregates the parameters based on the number of data samples in a client’s partition. The objective of the federated DiffTraj model is to minimise the loss of the model over the cross-client datasets, and can be formalised as
[…]
where $|X_{c}|$ and $|X_{\textnormal{glob}}|$ denote the number of trips in the dataset of client $c$ and in the global set of all clients, respectively. The process of training a federated DiffTraj model is shown in Algorithm 1. After the global model has been trained, each client is queried for conditional information which is then used to generate the central synthetic dataset, $X_{\textnormal{synth}}$. This conditional information can be centralised, since it does not compromise the privacy of the real trajectory data.
Algorithm 1 Training the DiffTraj (Zhu et al., 2023) model through FL using FedAvg.
@@ 4.1.1. Federated Framework
12:         Global updates $\theta_{r_{\textnormal{glob}}+1}\leftarrow\sum_{c=1}^{C}\frac{|X_{c}|}{|X_{\textnormal{glob}}|}\theta_{r_{\textnormal{glob}}}^{c}$
@@
To develop $X_{\textnormal{synth}}$, once the federated DiffTraj model is trained, client’s can be queried for conditional observations, proportionally based on the amount of orders in their dataset. This will ensure the conditional observations randomly sampled are evenly distributed for the range of clients. $|X_{c}|$ is available to the server in a FedAvg framework as shown in Equation (6). Synthetic data is only generated for a duration of 10 days, so that the ratio of synthetic regional inflow data to real regional inflow data for each client is 1:3.
4.2. Federated Traffic Flow Prediction Model
Using each client’s local individual trajectory dataset $X_{c}$, the regional inflow dataset, $\bm{Q_{c}}$, can be constructed, where $\bm{Q_{c}}\in\mathbb{R}^{TxN}$, $T$ is the number of regional traffic flow interval samples, and $N$ is the number of regions captured in the data. Each client can augment their local dataset with the synthetic regional inflow dataset $Q_{\textnormal{synth}}$ derived from the synthetic dataset $X_{\textnormal{synth}}$ shared by the server. Figure 3 illustrates the structure of the FedTPS framework for the traffic prediction task.
Figure 3. FedTPS framework for training a traffic prediction model in a federated setting, with synthetic data augmentation of each client’s local dataset.
[…]
System level diagram of server and clients in FL setting

### v2

@@ Abstract
Deep-learning based traffic prediction models require vast amounts of data to learn embedded spatial and temporal dependencies. The inherent privacy and commercial sensitivity of such data has encouraged a shift towards decentralized data-driven methods, such as Federated Learning (FL). Under a traditional Machine Learning paradigm, traffic flow prediction models can capture spatial and temporal relationships within centralised data. In reality, traffic data is likely distributed across separate data silos owned by multiple stakeholders. In this work, a cross-silo FL setting is motivated to facilitate stakeholder collaboration for optimal traffic flow prediction applications. This work introduces an FL framework, referred to as FedTPS, to augment each client’s local dataset with generated synthetic data by training a diffusion-based traffic flow generation model through FL. The proposed framework is evaluated on two large scale real world ride-sharing datasets and assessed against various generative and spatio-temporal prediction models, including a novel prediction model we introduce which leverages Temporal and Graph Attention mechanisms to learn embedded Spatio-Temporal dependencies. Experimental results show that FedTPS outperforms several FL baselines with respect to global model performance.
Index Terms:
@@
Federated Learning, Traffic Flow Prediction, Spatio-Temporal Modelling, Generative Models, Synthetic Data, Data Augmentation.
I Introduction
@@
Modern Intelligent Transport Systems (ITS) have increasing levels of connectivity enabling data exchange between users and infrastructure; it is projected that by 2026, all vehicles on UK roads will have some degree of connectivity [1]. This data can be leveraged to develop traffic prediction models, which can enhance the safety and productivity of users, as well as deliver significant social, environmental, and economic benefits to a city’s inhabitants. Deep learning-based models have demonstrated exceptional accuracy in their ability to model spatio-temporal relationships, and predict future traffic states [2, 3, 4, 5, 6].
@@
The majority of previous works towards spatio-temporal prediction aim to train a model from a centralised perspective where the collective data is available on a single device. Centralising city-wide ITS data in real-world settings is a difficult task; ITS represent a partnership between data-owning stakeholders such as local government authorities, ride-sharing companies, and micro-mobility organisations. Sharing journey data has implications on user privacy (GDPR), as well as commercial implications. The development of frameworks for decentralised data-driven approaches, which eliminate the need to centralize private and commercially sensitive information, is therefore critically important within the ITS sector for fostering stakeholder collaboration.
Figure 1: Federated framework for the traffic flow prediction task within an ITS. Each organisation collects data from their respective vehicle fleet, and through FL, collaboratively train a model to estimate traffic flow between regions (shown as orange grid cells).
Federated Learning (FL) [7] provides a solution for training traffic prediction models collaboratively by decentralising the training process of a traffic prediction model. Such a framework is detailed further in Figure 1. Although recent works in FL demonstrate its capacity to train a global model which learns spatio-temporal dependencies across distributed data, these works do not directly address the issue of data heterogeneity [8], or instead focus on the client selection strategy or parameter aggregation function [9, 10, 11, 12, 13, 14, 15].
To tackle the issue of data heterogeneity and imbalanced data partitions across clients, data sharing mechanisms for FL applications in highly heterogeneous settings were explored in [16] wherein a small partition of real data is distributed to clients to augment their local datasets. However, such an approach which shares real data between clients may violate privacy requirements.
@@
Research within FL for image classification tasks have recently looked at training generative models to augment local client datasets without sharing data that would compromise data privacy [17, 18, 19]. For applications beyond image classification where data contains spatial and temporal dependencies and heterogeneity is principally due to quantity and feature-based skews, limited work exists on the benefits of augmenting FL client datasets using synthetic data.
@@
Emerging research in the field of generative models have demonstrated the exceptional capacity for models to generate synthetic data that capture spatio-temporal interdependencies of traffic data [20]. Leveraging recent works in FL as well as in the field of generative models for spatio-temporal data generation, we present a framework for Federated Traffic Prediction with Synthetic data augmentation (FedTPS). This framework first trains a federated diffusion-based model for traffic flow data generation and generates a synthetic dataset developed from the learned global data distribution. A traffic flow prediction model is then trained through FL, wherein each client’s local dataset has been augmented with a generated synthetic dataset.
@@
In real-world implementations of traffic prediction models, the data owned by separate clients may contain inherent skews, contributing to data heterogeneity across client datasets. The proposed FedTPS framework directly inhibits the negative effects of data heterogeneity when training a global model through FL, while also increasing the amount of data in each client’s local dataset. To the best of our knowledge, this work is the first to explore the benefits of augmenting the federated framework with synthetic data in traffic prediction applications.
@@
  • We present a novel traffic prediction model, the Graph Attention and Temporal Attention Unit (GATAU), which builds upon the existing spatio-temporal model, the Temporal Attention Unit (TAU) [5], to also leverage Graph Attention mechanisms.
@@
  • We propose a novel FL framework, FedTPS, which trains a federated generative model to produce a synthetic traffic flow dataset. This synthetic data is used to augment client datasets for traffic flow prediction, and improve the predictive performance of the global FL model.
@@
  • We assess the performance of several traffic flow prediction models trained under an FL setting. Our experiments are validated on two real-world datasets to analyse the impact of data augmentation on traffic flow prediction model performance, and demonstrate improved results relative to various FL frameworks designed to address data heterogeneity across data-silos.
II Related Work
II-A Federated Learning
@@
The distributed nature of the data in an FL setting has a detrimental effect on the accuracy of the model relative to centralized training paradigms due to the fact that the distribution of data among clients is non-Independent and Identically Distributed (non-IID). In cross-silo FL settings, each client may represent a data center, owned by a particular organisation, which stores large amounts of data aggregated from many devices. FL has been extended to optimise the training of the global model [21], as well as to improve its applicability in non-IID settings [22]. Yet such methods focus on optimising the FL framework rather than address the data heterogeneity issues directly.
@@
Existing data augmentation approaches for FL focus on image classification tasks, and typically rely on Generative Adversarial Networks (GANs) [18, 17, 19], or on sharing mean latent representations of client data [23, 24]. Unlike image generation applications, traffic data generation requires modelling the dynamics of data over both spatial and temporal dimensions. Differentially-private representations of spatio-temporal data can be produced by altering real data [25, 26], or mixing partial aspects of the data [27]. These processes affect the spatio-temporal characteristics of the data itself, however. Recent work into diffusion-based models for spatio-temporal data generation demonstrate their capacity to generate high-fidelity synthetic trajectories which capture the spatio-temporal characteristics of sample traffic data, while ensuring the privacy of the real data.
@@
Existing work on FL applied for traffic prediction implement novel client selection strategies to cluster clients with similar spatial characteristics [14, 13, 9]. Other such works look to adapt the aggregation procedure of the FL setting [12, 10, 15]. FedTDP [8] explores the data heterogeneity issue which may be found in real-world traffic data silos, yet does not propose a method for dealing with the inherent data heterogeneity across client data silos.
II-B Traffic Flow Modelling
To extract temporal dependencies, deep-learning based models typically employ either Recurrent Neural Networks (RNNs) based architectures [3, 11], temporal Convolutional Neural Networks (CNNs) which convolve around the time dimension of traffic flow data [2, 4], or temporal attention mechanisms to capture the evolution of the data over time [5].
Spatial relationships embedded within data can be captured through CNNs which convolve around the spatial dimension of data [5]. Traffic data can also be represented as graph networks, enabling Graph Convolutional Networks (GCNs) to capture spatial relationships within the data. GCN operations are dependent on the graph structure, which requires expert contextual knowledge to capture inter-region relationships. Graph Attention Networks (GATs) [28] implement a self-attention mechanism to learn the relationship between nodes.
To capture spatio-temporal dependencies within data, models such as STGCN [2] implement GCNs with temporal CNNs to capture spatio-temporal dependencies within data, while DCRNN [3] leverages GCNs with RNN-based models. GraphWaveNet [4] also utilises temporal CNNs and GCNs, however it learns an adjacency matrix for the graph structured data. The Temporal Attention Unit (TAU) [5] utilises CNNs and temporal Attention mechanisms to capture both spatial and temporal relationships within data.
Generative approaches have been explored for synthesising realistic and diverse spatio-temporal data [29, 30]. Conditional Generative Adversarial Networks (cGANs) [31] extend the standard GAN architecture by conditioning both the generator and discriminator on auxiliary variables (e.g. traffic density, or environmental context) to enhance control over the generation process. Similarly, Conditional Variational Autoencoders (cVAEs) [32] integrate conditional information into the latent space of the VAE framework. DiffTraj [20] is a conditional diffusion-based trajectory generation model which leverages Traj-UNet, a network based on the UNet architecture [33], to model the noise transitions at each time step in the diffusion process. The model incorporates conditional information, embedded by a Wide and Deep network [34], to enhance the generative ability of the model.
III Problem Definition
Definition 1 (Traffic Flow Generation): To emulate the setting of various organisations each with ownership over separate data silos, the global trajectory dataset, $X_{\textnormal{glob}}$, is partitioned into $C$ clients such that $X_{\textnormal{glob}}=(X_{1}\cup X_{2}\cup...\cup X_{C})$, where $X_{c}$ is client $c$’s trajectory dataset. Traffic flow data is obtained from $X_{c}$ by discretising the city into $N$ regions, and defining $q_{c}^{t}\in\mathbb{R}^{N}$ as client $c$’s traffic inflow over $N$ regions at time step $t$. The objective of data generation is, given a set of traffic inflows, $Q_{c}=\{q_{c}^{1},q_{c}^{2},\ldots,q_{c}^{m}\}$, learn a model, $G$, which can generate a set of synthetic inflows, $Q_{\textnormal{synth}}=\{q_{\textnormal{synth}}^{1},q_{\textnormal{synth}}^{2},\ldots,q_{\textnormal{synth}}^{m}\}$, where the synthetic data retains the real data’s spatial-temporal characteristics. The federated task is to train model G over $X_{\textnormal{glob}}$ through FL using the distributed client datasets.
Definition 2 (Traffic Flow Prediction): Given historical traffic flow $Q=(q_{1},q_{2},...,q_{T})\in\mathbb{R}^{N\times T}$, train a model to predict the traffic inflow of all regions at future time step $T+\tau$, hence predict $q_{T+\tau}$. Based on the partitioned trajectory data $X_{\textnormal{glob}}$, client $c$ can derive its regional traffic inflow data $Q_{c}$, and can also augment its own local dataset using the synthetic dataset $Q_{\textnormal{synth}}$. The federated traffic flow prediction task is then to collaboratively train a global prediction model using the distributed client datasets with synthetic data augmentation, through FL.
IV Methodology
This work aims to address a real-world setting of multiple organisations, each with their independent dataset of vehicle trajectories, collaborating through FL to train a trajectory generation model and subsequent traffic flow prediction model. To replicate this scenario, a large-scale, real-world trajectory dataset is segmented into subsets representing local client datasets in a cross-silo FL setting. Within this FL setting, a server acts as a secure, trusted central system which coordinates federated training between clients.
The proposed FedTPS framework first aims to train a generative model through the partitioned vehicle trajectory datasets to generate a synthetic dataset from the global distribution of data. The second step aims to train a traffic flow prediction model using each client’s traffic flow data, and augment each client’s data with the synthetic generated data. This will reduce the effects of data heterogeneity present across the client datasets, and increase the amount of training data available to clients.
IV-A Federated Generative Model
As a probabilistic generative model, a diffusion model aims to generate synthetic data by modeling two sequential processes: 1) A series of forward processes which gradually injects noise into the data, and 2) The reverse of these processes, which learn to capture the original data distribution from a noisier version [35]. Recent works have applied diffusion-based models for high-fidelity data generation in various domains [36, 37, 38]. The forward process is defined as a Markov chain with $F$ Gaussian transitions that map the real data $x_{0}\sim g(x_{0})$ to $x_{F}$, a latent variable with the same dimensionality as $x_{0}$, according to
$g(x_{1},...,x_{F}|x_{0})=\prod^{F}_{f=1}g(x_{f}|x_{f-1}),$ (1)
$g(x_{f}|x_{f-1})=\mathcal{N}(x_{f};\sqrt{1-\beta_{f}}x_{f-1},\beta_{f}\mathbf{I}),$ (2)
$p_{\theta}(x_{0},...,x_{f-1}|x_{F})=p(x_{F})\prod^{F}_{f=1}p_{\theta}(x_{f-1}|x_{f}),$ (3)
$p_{\theta}(x_{f-1}|x_{f})=\mathcal{N}(x_{f-1};\mu_{\theta}(x_{f},f),\sigma_{\theta}(x_{f},f)^{2}\mathbf{I}),$ (4)
$\mathcal{L}_{c}(\theta):=\mathbb{E}_{x_{0}\in X_{c},\epsilon,f}\left[||\epsilon-\epsilon_{\theta}(x_{f},f)||^{2}\right].$ (5)
The UNet architecture [33] has proved highly effective in generative modeling tasks; the fusion of multiple scales allows a diffusion model built on UNet to robustly navigate noise removal in each denoising step of the diffusion process. UNet employs a contracting encoder, which applies convolutional operations to progressively reduce spatial resolution and extract salient features, followed by an expanding decoder that upsamples these encoded representations to restore the original spatial dimensions. Skip connections link corresponding encoder and decoder layers to retain spatial information from multiple scales that may otherwise be lost during downsampling; as the decoder upsamples the encoded input, the skip connections provide finer-scale information from the earlier encoded layers.
For the trajectory generation task, Zhu et al. [20] introduce an attention-driven transition module, and demonstrate its ability to bolster the spatio-temporal modeling capacity of diffusion-based methods. Akin to their DiffTraj model, our Traffic Flow Diffusion model (TFDiff) utilises the UNet architecture, with an intermediary attention block.
Conditional diffusion models enable additional information to be embedded into the reverse denoising process as conditional variables into Eq. (4). TFDiff is conditioned on information regarding the time interval, the day of the week, the maximum traffic flow at any region within the time interval, and the region index with the maximum traffic flow.
IV-A1 Federated Framework
In a scenario where the global trajectory dataset is segmented across data silos from multiple organisations, a generative model cannot be trained under a traditional framework since the data is not centralised.
Figure 2: FedTPS Framework for training a federated generative model and subsequent traffic flow prediction model. The first stage, shown as the purple dashed line, trains the federated diffusion model to develop the FedTPS synthetic dataset. The second stage, shown as the orange dashed line, trains the federated traffic flow prediction model with the client’s local dataset, as well as the synthetic data. $X^{\textnormal{CO}}_{c}$ represents client $c$’s conditional observations required for the TFDiff model.
Under a federated setting, the updated parameters from each client’s local training on the global model are communicated to a central server which then aggregates the results to update the global model. FedAvg aggregates the parameters based on the number of data samples in a client’s partition. The objective of the federated TFDiff model is to minimise the loss of the model over the cross-client datasets, and can be formalised as
@@ IV-A1 Federated Framework
where $|Q_{c}|$ and $|Q_{\textnormal{glob}}|$ denote the number of traffic flows in the dataset of client $c$ and in the global set of all clients, respectively. The process of training a federated Diff model is illustrated in step one of Figure 2, and is detailed further in Algorithm 1. Once the global TFDiff model has been trained, it is shared with all clients, who then use it to generate the synthetic datasets to augment their respective local datasets.
Algorithm 1 Training the TFDiff model through FL using FedAvg.
@@
12:         Global updates $\theta_{r_{\textnormal{glob}}+1}\leftarrow\sum_{c=1}^{C}\frac{|Q_{c}|}{|Q_{\textnormal{glob}}|}\theta_{r_{\textnormal{glob}}}^{c}$
@@
The conditional data derived from each client’s private dataset, as well as the distributed global TFDiff model, are used by each client to generate their synthetic dataset. Synthetic data is generated for a duration of 10 days. Generating additional synthetic data did not yield significant improvement in the global federated traffic prediction model.
IV-B Federated Traffic Flow Prediction Task
Using each client’s local individual trajectory dataset $X_{c}$, the regional inflow dataset, $Q_{c}$, can be constructed, where $Q_{c}\in\mathbb{R}^{T\times N}$, $T$ is the number of regional traffic flow interval samples, and $N$ is the number of regions within the city. Each client can augment their local dataset with the synthetic regional inflow dataset $Q_{\textnormal{synth}}$ generated from each client using the TFDiff model for inference. Step two in Figure 2 illustrates the structure of the FedTPS framework for the traffic prediction task.